# Variational Self-Supervised Learning

### v1

@@ III-A Variational Formulation
$\mathcal{L}(x)=\mathbb{E}_{q(z\mid x)}[\log p(x\mid z)]-D_{\mathrm{KL}}(q(z\mid x)\|p(z))$
@@ III-C EMA-Updated Teacher Network as Prior
$\theta_{t}\leftarrow\tau\theta_{t}+(1-\tau)\theta_{s}$
@@
$p_{\theta_{t}}(z\mid x_{1})=\mathcal{N}(z\mid\mu_{t}(x_{1}),\sigma_{t}^{2}(x_{1}))$
@@
$D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
@@ III-D Reconstruction Term with Denoiser Networks
$p_{\phi}(z\mid\mu_{1},\sigma_{1})=\mathcal{N}(z\mid\mu_{\phi}(x_{1}),\sigma_{\phi}^{2}(x_{1}))$
@@
$\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
@@
  • The student encoder, $q_{\theta_{s}}(z\mid x_{2})$, to map $x_{2}$ to latent space.
@@ III-E Derivation of the Self-Supervised ELBO
We derive the modified ELBO for this self-supervised framework. Given views $x_{1}$ and $x_{2}$, we maximize the likelihood $\log p_{\phi}(z\mid\mu_{1},\sigma_{1})$. Applying Jensen’s inequality gives:
$\log p_{\phi}(z\mid\mu_{1},\sigma_{1})$ $\geq\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad-D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
Thus, the modified ELBO becomes:
$\mathcal{L}(x_{1},x_{2})$ $=\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad-D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
Likelihood Terms: Encourage consistency between $x_{2}$’s latent samples and $x_{1}$’s predicted distribution. KL Term: Regularizes the student’s posterior to align with the EMA-stabilized teacher prior.
III-F Implementation Details
@@
Variational Modeling. Latent variables are modeled as Gaussian distributions with parameters constrained via the Softplus activation on variance terms. Stability is enhanced by adding noise proportional to $0.01\cdot\epsilon^{2}$, where $\epsilon\sim\mathcal{N}(0,1)$.
@@
Cosine-Based KL and Log-Likelihood. To enhance semantic alignment in high-dimensional latent spaces, we propose cosine-based alternatives to classical KL divergence and log-likelihood formulations.
@@
$\tilde{D}_{\mathrm{KL}}^{\cos}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})$ $=\frac{1}{2}\Big{[}\log(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+(1-\cos(\mu_{1},\mu_{2}))^{2}$
$\quad+(1-\cos(\sigma_{1},\sigma_{2}))-1\Big{]}$
@@
~$\tilde{\mathcal{L}}^{\cos}_{\text{ll}}(\mu_{1},\mu_{2},$ $\sigma_{1}^{2},\sigma_{2}^{2})=\log(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+4(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+(1-\cos(\mu_{1},\mu_{2}))^{2}(1-\cos(\sigma_{1},\sigma_{2}))$
These formulations leverage:
  • Log-term: captures similarity in uncertainty.
  • Variance contrast: emphasizes alignment of distributional spread.
  • Mean similarity: enforces semantic closeness, weighted by variance agreement.
Cosine similarity provides a robust angular measure, particularly effective in high-dimensional embedding spaces, offering an alternative to Euclidean-based alignment in variational settings.
@@
17:        Compute total loss: $\mathcal{L}=\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{ll}}$
@@ IV Experimental Evaluation
VSSL (this work) 93.08 70.53 81.50 81.34
@@ V Results and Discussion
Table I presents the top-1 classification accuracy (%) of several prominent self-supervised learning (SSL) methods across three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet-100. We evaluate each method under both online and offline linear evaluation protocols.
Our proposed method, VSSL, consistently demonstrates strong performance across all benchmarks. On CIFAR-10, VSSL achieves an accuracy of 93.08%, ranking second only to MoCo V3 (93.10%) by a narrow margin of 0.02%. On CIFAR-100, VSSL achieves 70.53%, placing it among the top three performers, slightly behind Barlow Twins (70.90%) and BYOL (70.46%).
On the more challenging ImageNet-100 dataset, VSSL outperforms all competing methods under both evaluation protocols. Specifically, it achieves 81.50% in the online setting and 81.34% in the offline setting, surpassing the closest competitors, BYOL and MoCo V3, by over 1%.
These results highlight the effectiveness of our approach in capturing transferable visual representations. VSSL not only maintains competitive performance on relatively simpler datasets (e.g., CIFAR-10/100) but also excels on large-scale, diverse datasets such as ImageNet-100, where robust representation learning is crucial.
Interestingly, methods like Barlow Twins and BYOL show strong results on CIFAR datasets but exhibit diminished performance on ImageNet-100. In contrast, approaches such as MoCo V3 and NNCLR achieve more balanced results across different data scales. VSSL appears to combine the strengths of both, suggesting better generalization and robustness.
Overall, the consistent superiority of VSSL across different datasets and evaluation settings underscores its potential as a versatile and scalable SSL framework. Further ablation studies and transfer learning experiments may provide additional insight into the key components contributing to its performance.
@@ Acknowledgment
The numerical calculations reported in this paper were fully/partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).

### v2

@@ III-A Variational Formulation
$\mathcal{L}(x)=\mathbb{E}_{q(z\mid x)}[\log p(x\mid z)]-D_{\mathrm{KL}}(q(z\mid x)\|p(z))$ (1)
@@ III-C EMA-Updated Teacher Network as Prior
$\theta_{t}\leftarrow\tau\theta_{t}+(1-\tau)\theta_{s}$ (2)
@@
$p_{\theta_{t}}(z\mid x_{1})=\mathcal{N}(z\mid\mu_{t}(x_{1}),\sigma_{t}^{2}(x_{1}))$ (3)
@@
$D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$ (4)
@@ III-D Reconstruction Term with Denoiser Networks
$p_{\phi}(z\mid\mu_{1},\sigma_{1})=\mathcal{N}(z\mid\mu_{\phi}(x_{1}),\sigma_{\phi}^{2}(x_{1}))$ (5)
@@
$\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$ (6)
@@
  • The teacher encoder, $q_{\theta_{t}}(z\mid x_{2})$, to map $x_{2}$ to latent space.
@@ III-E Derivation of the Self-Supervised ELBO
We derive the modified ELBO for this self-supervised framework. Given views $x_{1}$ and $x_{2}$, we maximize the likelihood $\log p_{\phi,z\sim q_{\theta_{t}}(z\mid x_{2})}(z\mid\mu_{1},\sigma_{1})$. Applying Jensen’s inequality gives:
$\log p_{\phi,z\sim q_{\theta_{t}}(z\mid x_{2})}(z\mid\mu_{1},\sigma_{1})\geq$
$\qquad+\mathbb{E}_{z\sim q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\qquad-D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$ (7)
III-F Overall Objective
Putting it all together, our variational self-supervised loss is
$\mathcal{L}_{\mathrm{VSSL}}$ $=\sum_{v_{1}=1}^{2}\sum_{\begin{subarray}{c}v_{2}=1\end{subarray}}^{2}\Bigl{[}D_{\mathrm{KL}}\bigl{(}q_{\theta_{s}}(z\mid x_{v_{1}})\,\|\,p_{\theta_{t}}(z\mid x_{v_{2}})\bigr{)}$
$\quad-\;\mathbb{E}_{z\sim q_{\theta_{s}}(z\mid x_{v_{1}})}\bigl{[}\log p_{\phi}(z\mid\mu_{v_{2}},\sigma_{v_{2}})\bigr{]}\Bigr{]},$ (8)
and we solve
$\min_{\theta_{s},\phi}\;\mathcal{L}_{\mathrm{VSSL}}.$
III-G Implementation Details
@@
Variational Modeling. Latent variables are modeled as Gaussian distributions with unconstrained mean and variance parameters and an approximation to actual re-parametrization trick:
$\epsilon\sim\left|\mathcal{N}(0,1)\right|,\quad z=\mu+\sigma\cdot\epsilon$
@@
Cosine-Based KL Divergence and Log-Likelihood. To enhance semantic alignment in high-dimensional latent spaces, we propose cosine-based alternatives to classical KL divergence and log-likelihood formulations, leveraging the angular properties of cosine similarity.
@@
$\tilde{D}_{\mathrm{KL}}$ (${}^{\cos}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\frac{1}{2}\Big{[}\log\big{(}1-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+\big{(}\mathrm{softplus}_{\beta=3}\big{(}-\cos(\mu_{1},\mu_{2})\big{)}\big{)}^{2}$
$+\big{(}\mathrm{softplus}_{\beta=3}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}-1\big{)}\Big{]}$
@@
~$\tilde{\mathcal{L}}^{\cos}_{\mathrm{ll}}$ $(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\log\big{(}1-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+4\,\mathrm{softplus}_{\beta=1}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+\Big{[}\big{(}\mathrm{softplus}_{\beta=1}\big{(}-\cos(\mu_{1},\mu_{2})\big{)}\big{)}^{2}$
×$\qquad\times\mathrm{softplus}_{\beta=1}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}\Big{]}$
The softplus function with temperature parameter $\beta$ is defined as:
$\mathrm{softplus}_{\beta}(x)=\frac{1}{\beta}\log\big{(}1+\exp(\beta x)\big{)}$
Cosine similarity provides a robust angular measure, particularly effective in high-dimensional embedding spaces, offering an alternative to Euclidean-based alignment in variational frameworks.
@@
17:        Compute total loss: $\mathcal{L}=\mathcal{L}_{\text{KL}}-\mathcal{L}_{\text{ll}}$
@@ IV Experimental Evaluation
VSSL (this work) 93.35 70.75 81.15 81.01
@@ V Results and Discussion
Table I presents the top-1 classification accuracy (%) for a selection of prominent self-supervised learning (SSL) methods across three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet-100. Evaluations are conducted under both online and offline linear evaluation protocols to ensure a comprehensive assessment of learned representations.
For CIFAR-10, the majority of methods achieve over 90% accuracy, demonstrating the maturity of SSL techniques on this relatively simple dataset. Notably, our proposed method, VSSL, achieves the highest top-1 accuracy at 93.35%, outperforming the strong baseline MoCo V3 (93.10%) and BYOL (92.58%).
On the more challenging CIFAR-100 dataset, which contains a higher number of classes and greater inter-class variability, Barlow Twins achieves the highest accuracy of 70.90%. VSSL closely follows with 70.75%, marking it as highly competitive and outperforming other methods such as MoCo V2+ (69.89%) and VICReg (68.54%). This demonstrates the effectiveness of VSSL in handling fine-grained classification tasks.
For the ImageNet-100 dataset, which provides a larger-scale evaluation, results are reported under both online and offline linear probing protocols. VSSL consistently outperforms all competing methods, achieving top-1 accuracies of 81.15% (online) and 81.01% (offline). These results surpass the second-best method, MoCo V3, which achieves 80.36% in both settings. The strong performance of VSSL across both evaluation modes indicates the robustness and generalizability of its learned representations.
Overall, the results suggest that VSSL not only achieves state-of-the-art performance across multiple datasets but also demonstrates superior generalization across diverse evaluation protocols. Its competitive accuracy on CIFAR-100 and ImageNet-100, which feature more complex classification scenarios, highlights the efficacy of variational self-supervised learning strategies. The consistent outperformance of established methods such as MoCo V3, BYOL, and Barlow Twins further underscores the potential of VSSL as a versatile and scalable SSL framework.
@@ Acknowledgment
The numerical calculations reported in this paper were fully performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).